OpenStudy (anonymous):

What is the solution to the equation 9 ^3x ≈ 7?

OpenStudy (anonymous):

Is it \[9^{3x}\approx 7\] ?

OpenStudy (anonymous):

yes

OpenStudy (dls):

\[\Huge 9^{3x}\approx 7\] Let us assume it nearly equal. \[\Huge 3x = \log_9 7\] \[\Huge x= \frac{\log_9 7}{3}\]

OpenStudy (dls):

= \[\Huge \frac{\log 7}{3 \log 9 } = \frac{\log 7}{6 \log 3} \]

OpenStudy (dls):

x is nearly \(\Huge 0.30\)

OpenStudy (anonymous):

here are the answer choices x = 0.376 x = 0.295 x = −0.295 x = −0.376

OpenStudy (dls):

You can verify it by \[\Large 9^{3 \times 0.30} = 9 ^{ 0.9 } =7.22 ..\] nearly 7 x is also NEARLY 0.30 (round off..)

OpenStudy (dls):

0.295 to be exact which I rounded upto 0.30

OpenStudy (anonymous):

thank you

OpenStudy (dls):

uw